jdishdjaujssh:

Leigh bought a bag of beads 1/3 of the beads were green, 1/9 were black and 1/5 of the remainder were white. If there were 25 white beads how many beads were there altogether

AZ:

We have a bag of beads 1/3 of the beads are green 1/9 are black of the REMAINING beads, 1/5th of those beads are white so first we need to figure out how many beads are remaining that 1/5 of them were white so what is 1 - 1/3 - 1/9 = ??

AZ:

and so whatever fraction you get, multiply that with 1/5 this new fraction you get will tell you how many beads are white out of the total number of beads so what you have to do to get your final answer: let's say, you get a/b is the fraction of white marbles to total you have to divide that from the number of white marbles there and that will give you the total number of marbles so \( 25 \div \dfrac{a}{b}\) and that is going to be equal to \( 25\times \dfrac{b}{a}\)